During the scheduled downtime in August the ZnS lenses in the fore-optics and spectrograph were removed together with the ZnSe prism in the spectrograph. These components were then returned to Optics For Research for removal of the anti-reflection coats before being sent to Thin Film Labs for recoating to fix the low throughput in the 0.8-0.9 micron range. SpeX was reassembled and cooled down in time for the first scheduled run on August 30. Unfortunately the rotator and order sorter mechanisms failed and the instrument was immediately warmed up, resulting in the loss of one three-night run (Vacca). Fire brigade action by the staff fixed two loose couplings and had the instrument operational for the next observer (Binzel Sep 2). SpeX worked smoothly throughout the remainder of the quarter.
One area of relatively slow progress has been the writing of the IDL-based spectral extraction code. The main task of the spectral extraction tool is to remove the instrument signature from the data and to provide observers with a tool to assess signal-to-noise whilst observing. This involves automated flatfielding, spectral extraction and wavelength calibration, in a user-friendly GUI. The coding has proved to be more complicated than anticipated and this has slowed the reduction of spectral data. A beta version was made available to observers during this quarter.